40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Jul 16 23:28:24 2021, Koji, Summary, Cryo vacuum chamber, Sus Test Work 07/16/2021 7x
    Reply  Sat Jul 17 00:55:41 2021, Koji, Summary, Cryo vacuum chamber, Temp Log 210716_2255 temp_log_cool_down_20210716_2255.pdf
       Reply  Mon Jul 19 15:57:17 2021, Stephen, Summary, Cryo vacuum chamber, Temp Log 210716_2255 
          Reply  Mon Jul 19 17:21:19 2021, Koji, Summary, Cryo vacuum chamber, Temp Log 210716_2255 temp_log_cool_down_20210716_2255.pdf
             Reply  Tue Jul 20 11:33:52 2021, Koji, Summary, Cryo vacuum chamber, A cooling model (Temp Log 210716_2255) cooling_model.pdfIMG_9390.JPGIMG_9391.JPG
                Reply  Tue Jul 20 17:28:30 2021, Koji, Summary, Cryo vacuum chamber, A cooling model (Temp Log 210716_2255) cooling_model.pdf
                   Reply  Wed Jul 21 21:05:59 2021, Koji, Summary, Cryo vacuum chamber, Test mass cooling (2021/07/16 ~ 2021/07/21) oplev_trend.png20210721201333_IMG_0765.jpeg20210716234113_IMG_0742.jpegScreenshot_from_2021-07-21_20-19-09.png
                      Reply  Thu Jul 22 22:03:45 2021, Koji, Summary, Cryo vacuum chamber, Test mass heating in progress (2021/07/21 ~ 2021/07/23) IMG-9395.jpgtemp_log_warmup_20210721_2052.pdf
                      Reply  Sun Jul 25 21:45:46 2021, Koji, Summary, Cryo vacuum chamber, About the radiation heat transfer model 
Message ID: 2605     Entry time: Fri Jul 16 23:28:24 2021     In reply to: 0     Reply to this: 2606
Author: Koji 
Type: Summary 
Category: Cryo vacuum chamber 
Subject: Sus Test Work 07/16/2021 

[Stephen Koji]

We started cooling down of the test mass.

Venting

- Stephen vented the chamber at 2PM. An optical port was moved to see the OSEM from the back.

OSEM wiring

- Brought DSub crimp sockets from the 40m. We picked up 3x 1m LakeShore WCT-RB-34-50 (twisted silver-plated copper, 34 AWG with Teflon insulation). The ends of the wires were dangled so that crimping is possible. A single wire resistance was measured to be ~1Ohm at room temp. (Attachment 1)

- OSEM pin out / backside view (cable going down) (Attachment 2)

|   o   o   o |
| o   o   o   |                 Wire
  ^ ^ ^ ^ ^ ^---PD K        ---- R3
  | | | | |-----PD A        ---- B3
  | | | |-------LED A       ---- B2
  | | |---------LED K       ---- R2
  | |-----------Coil End    ---- B1
  |-------------Coil Start  ---- R1


Twisted Pair 1: (R1&B1) with 1 knot  at the feedthru side
Twisted Pair 2: (R2&B2) with 1 knot  at the feedthru side
Twisted Pair 3: (R3&B3) with 1 knot  at the feedthru side

Dsub feedthru in-air pinout (Mating side)

    1  2  3  4  5
\ o  o  o  o  o /
 \ o  o  o  o  /
   6  7  8  9

Pin1 - Coil Start
Pin6 - Coil End
Pin2 - LED K
Pin7 - LED A
Pin3 - PD A
Pin8 - PD K

Pin1-6 R=16Ohm
Pin2-7 Diode V (with Fluke) 1.18V (Pin2 black probe / Pin7 red probe)
Pin3-8 Diode V (with Fluke) 0.7V (Pin3 red probe / Pin8 black probe)

- OSEM pin out / backside view (cable going down)

Suspension installation (Attachment 3)

- The sus frame was moved into the chamber

- We measured the test mass dimension before installation: L 3.977" D 4.054"

- The attached mirror size is 1"x1" made of SUS #8 (?)

- The mass was suspended. The height / rotation of the mass was adjusted so that the reflecting mirror is visible from the oplev window and also the OSEM magnet is visible from the OSEM window.

- The OSEM was placed on an improvised holder. (Attachment 4)

Oplev installation

- ...Just the usual oplev installation. Adjusted the alignment and the return beam hits right next to the laser aperture. This beam was picked off by a mirror and steered into a QPD. (Attachments 5/6)

- The lever arm length is ~38" (960mm) -- 9" internal / 29" external
- The oplev signal is shaking so much and occupying ~50% of the full scale. Added a lens with f=250 to make the beam bigger, but the improvement was limited.

Pumping down

- Started ~8:30PM?

DAQ setup

- Wired 3 BNC cables from the table to the DAQ rack. CHX/Y/S are connected to ADC16/1718ch.

- The real-time processes seemed dead. Looked at [QIL ELOG 2546] to bring them up. TIM/DAQ error remains, but the data stream seems alive now. Leave it as it is.

Cooling

- Temp Logging started. Filename: temp_log_cool_down_20210716_2255.txt

- Cryocooler turned on. ~10:55PM

- Confirmed the cold head temp was going down. The cold head temp is 75K at 0:30AM

OSEM photo

- An example photo was taken from the rear window. The attempt with 40m's Canon failed. Attachment 7 was taken with KA's personal compact camera with a smartphone LED torch. The gap between magnet and OSEM is highly dependent on the view axis. So this is just a reference for now.

Attachment 1: 20210716170727_IMG_0719.jpeg  1.270 MB  Uploaded Sat Jul 17 01:13:19 2021  | Hide | Hide all
20210716170727_IMG_0719.jpeg
Attachment 2: 20210716174712_IMG_0723.jpeg  2.223 MB  Uploaded Sat Jul 17 01:13:46 2021  | Hide | Hide all
20210716174712_IMG_0723.jpeg
Attachment 3: 20210716195953_IMG_0726.jpeg  3.320 MB  Uploaded Sat Jul 17 01:14:03 2021  | Hide | Hide all
20210716195953_IMG_0726.jpeg
Attachment 4: 20210716200005_IMG_0728.jpeg  1.987 MB  Uploaded Sat Jul 17 01:14:03 2021  | Hide | Hide all
20210716200005_IMG_0728.jpeg
Attachment 5: 20210716200224_IMG_0734.jpeg  3.884 MB  Uploaded Sat Jul 17 01:14:12 2021  | Hide | Hide all
20210716200224_IMG_0734.jpeg
Attachment 6: 20210716200112_IMG_0733.jpeg  1.520 MB  Uploaded Sat Jul 17 01:14:12 2021  | Hide | Hide all
20210716200112_IMG_0733.jpeg
Attachment 7: 20210716234113_IMG_0742.jpeg  754 kB  Uploaded Sat Jul 17 01:31:24 2021  | Hide | Hide all
20210716234113_IMG_0742.jpeg
ELOG V3.1.3-